OpenStudy (anonymous):

Solve the following equation. Then place the correct numbers in the boxes provided. (Write the answers as improper fractions, with positive answers first.) Because | 3x - 4 | = 15, x = or x =

OpenStudy (anonymous):

3x-4=15 or 3x-4=-15

OpenStudy (anonymous):

19/3>x>-11/3

OpenStudy (anonymous):

x=19/3 or x=-11/3